OpenStudy (anonymous):

Can someone help me out to Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.

OpenStudy (anonymous):

parallel lines have same slope

OpenStudy (anonymous):

so slope of the required line will be -1 too and the eqn will be: y-2=-1(x+4)

OpenStudy (anonymous):

And then I have to simplify that? to y-2=-1x+-4 but I dont know if it stays positive or it changes to negative?

OpenStudy (anonymous):

yes, it will be y-2=-x-4 y=-x-2